Examiner's comment
In (b)(i) the hypotheses were usually correctly stated with most students using a one-tailed test. On some occasions students are writing these hypotheses in words rather than using appropriate notation. Most students were able to select and use the correct Binomial distribution. Many, however, then calculated P(X = 11) or incorrectly attempted P(X >= 11) = 1 – P(X <= 11). Of the students who gained the A mark for the probability 0.16, the conclusion was usually correct and mostly given in sufficient contextual detail. Those students who used a critical region approach generally did less well, as most omitted the required p-value and scored a maximum of 3 marks. relevant probabilities at some point in part (b). A few students gave a probability rather than a region as their final answer here.